On January 2, 2015, Grouper Corporation issued $2,050,000 of 10% bonds at 96 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”)

The bonds are callable at 101 (i.e., at 101% of face amount), and on January 2, 2020, Grouper called $1,230,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Grouper as a result of retiring the $1,230,000 of bonds in 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Loss on redemption $ ????


Prepare the journal entry to record the redemption. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation   

Debit

Credit

January 2, 2020

?

?

?

?
Answer #1

Face value of bonds = 2050000

Discount on bonds payable = 2050000*4/100 = 82000

Carrying value of January 2020 = 2050000-41000 = 2009000

Carrying value of 1230000 face value bonds = 2009000*1230000/2050000 = 1205400

a) Loss on redemption = (1230000*1.01)-1205400 = 36900

b) Journal entry

Date account and explanation Debit Credit
Jan 2 Bonds payable 1230000
Loss on redemption 36900
Discount on bonds payable 24600
Cash 1242300